Williamson Industries has $3 billion in sales and $1.8 billion in fixed assets. Currently, the company's fixed assets are operating at 90% of capacity.

a. What level of sales could Williamson Industries have obtained if it had been operating at full capacity? Write out your answer completely. For example, 25 billion should be entered as 25,000,000,000. Round your answer to the nearest cent.
$____

b. What is Williamson's target fixed assets/sales ratio? Round your answer to two decimal places.
___ %

c. If Williamson's sales increase 15%, how large of an increase in fixed assets will the company need to meet its target fixed assets/sales ratio? Write out your answer completely. For example, 25 billion should be entered as 25,000,000,000. Round your answer to the nearest cent. Negative amount should be indicated by a minus sign. Do not round intermediate calculations.
$____
Answer #1

a. Let sales at full capacity be x.

So, x*0.90 = $3,000,000,000

x = $3,000,000,000/90 = $3,333,333,333.33

So, sales at full capacity is $3,333,333,333.33.

b. target fixed assets/sales ratio = fixed assets/sales

= $1,800,000,000/$3,000,000,000

= 0.6 or 60.00%

c. New sales after 15% increase = $3,000,000,000*1.15 =$3,450,000,000

Let the new fixed asset after 15% increase in sales be x.

target fixed assets/sales ratio = fixed assets/sales

0.60 = x/$3,450,000,000

x = $3,450,000,000*0.60 = $2,070,000,000

increase in fixed assets = $2,070,000,000 - $1,800,000,000 = $270,000,000